A total of $34000 was invested in two parts, one part at an interest of 20% and the other part at an interest of 15%. The amount invested at 20% invest was $500 more than the amount invested at 15% interest. Find the amount invested in each part.
Answer by checkley71(8403)

34000=x+(x+500)
34000=2x+500
2x=34000-500
2x=33500
x=33500/2
x=$16,750 amount invested @ 15%
34,000-16,750=$17,250 invested @ 20%